### v1

# CleverDistiller: Simple and Spatially Consistent Cross-modal Distillation

Abstract
Vision foundation models (VFMs) such as DINO have led to a paradigm shift in 2D camera-based perception towards extracting generalized features to support many downstream tasks. Recent works introduce self-supervised cross-modal knowledge distillation (KD) as a way to transfer these powerful generalization capabilities into 3D LiDAR-based models. However, they either rely on highly complex distillation losses, pseudo-semantic maps, or limit KD to features useful for semantic segmentation only. In this work, we propose CleverDistiller, a self-supervised, cross-modal 2D-to-3D KD framework introducing a set of simple yet effective design choices: Unlike contrastive approaches relying on complex loss design choices, our method employs a direct feature similarity loss in combination with a multi layer perceptron (MLP) projection head to allow the 3D network to learn complex semantic dependencies throughout the projection. Crucially, our approach does not depend on pseudo-semantic maps, allowing for direct knowledge transfer from a VFM without explicit semantic supervision. Additionally, we introduce the auxiliary self-supervised spatial task of occupancy prediction to enhance the semantic knowledge, obtained from a VFM through KD, with 3D spatial reasoning capabilities. Experiments on standard autonomous driving benchmarks for 2D-to-3D KD demonstrate that CleverDistiller achieves state-of-the-art performance in both semantic segmentation and 3D object detection (3DOD) by up to 10% mIoU, especially when fine tuning on really low data amounts, showing the effectiveness of our simple yet powerful KD strategy.
Figure 1: We observe how CleverDistiller (top) improves over the baseline ScaLR (bottom). 3D representations from both KD methods are fine-tuned to the 3D semantic segmentation task on nuScenes. We observe how our approach encourages spatially consistent semantic outputs.
1 Introduction
For many vision applications such as autonomous driving, research is focused on improving generalized understanding of the environment to increase safety and autonomy. However, constant updates of sensor types and setups from different manufacturers makes it difficult to train one perception system that excels on all use cases. Most perception models for, e.g., object detection or semantic segmentation, are known for their lack of robustness towards weather conditions, lighting, sensor setup, or driving environments that usually deviate between inference and training. Especially, LiDAR-based 3D models are known for their adaptation challenges [70, 82, 85, 86, 59, 72, 71, 55, 41, 67, 47, 89, 34, 2, 26].
Due to these inherent difficulties in adaptation, recent focus has shifted towards using VFMs such as DINO [36, 54]. These models provide features that are robust to different sensor setups and environments, while being able to support multiple downstream tasks [73, 45, 12, 63]. While many works address and extend VFMs in camera-based setups, our work is aimed at transferring VFM knowledge to more challenging LiDAR-based 3D networks. In this direction, cross-modal knowledge distillation has been used to transfer learned feature representations between the backbones of different sensor modalities. Initial methods focused mainly on distilling LiDAR-based features into camera-based models [18, 38]. However, with the emergence of VFMs the reverse direction, to distill generalized semantic 2D features into 3D LiDAR-based networks, has become highly relevant [57, 49, 83, 84]. While initial work is mostly focused on 3D semantic segmentation as a downstream task, we aim to obtain a pre-trained LiDAR-based backbone that can be applied to various downstream tasks, e.g., both 3D semantic segmentation and 3DOD.
Many previous approaches for cross-modal KD have tried to explicitly model the semantic dependencies by using semantic priors (generated by e.g., Segment Anything (SAM) or OpenSeeD [37, 90, 92, 83, 84]) and auxiliary spatial tasks. Although showing improvements in performance and robustness, these approaches tend to complicate the cross-modal KD process. Therefore, we build on the second type of approach presented by ScaLR [57] which proposes a simple distillation solution, however, without achieving SOTA performance. As demonstrated in Fig. 1 ScaLR often yields semantic outputs that are not consistent with the scene’s spatial structure.
To address these shortcomings, we propose CleverDistiller, a fully self-supervised KD framework that transfers features from a camera-based VFM to a LiDAR-based 3D network. Our approach relies on a multi-task feature distillation strategy that integrates cross-modal KD from a VFM and an auxiliary spatial task, i.e., occupancy prediction [8], to capture both semantic and spatial information in a unified framework, thereby improving spatial and semantic feature consistency, cf. Fig. 1. Our contributions can be summarized as follows:
  • We identify linear projection of 3D LiDAR features into 2D space of camera-based VFM features as the main limitation of simple cross-modal KD loss functions. We therefore introduce an MLP projection head on top of the 3D LiDAR backbone to allow flexible learning of 3D spatial properties in the distillation between structurally very different 2D image and 3D LiDAR features while retaining simplicity on the distillation loss from VFM features.
  • We also incorporate the auxiliary self-supervised task of occupancy prediction to preserve spatial information within the LiDAR model. Thereby, generalized semantic features distilled from the VFM are further encouraged to be spatially consistent across the 3D scene.
  • We demonstrate that CleverDistiller outperforms previous approaches in standard 2D-to-3D KD benchmarks for both 3D semantic segmentation and 3DOD. Our model is particularly successful when fine-tuned on low ammount of data, highlighting the generalization capabilities of our learned 3D features.
2 Related work
Pre-training Vision Models: Supervised pre-training on large labeled datasets [21, 44, 31, 65, 66] has being replaced by self-supervised pre-training [25]. Self-supervised pre-training involves training a model using a pretext task on large unlabeled datasets. Currently, self-supervised learning (SSL) methods predominantly use a contrastive  [14] or reconstruction formulation [96, 54]. Initially, a contrastive objective was considered over data instances in a batch [78, 14] or a memory bank [32, 15, 16]. Recent works such as DINO [12] contrast over prototypical clusters [46, 10, 3, 11, 29] and learn representations that are highly data efficient [4, 58]. Reconstruction based methods [33, 6] typically lack data efficiency [5]. Then, DINOv2 [54] consolidated a range of strategies such as masked image modeling [96] and produced a model that excels on a wide range of downstream tasks, adopted in many recent works [30, 24, 68]. In this work we also adopt DINOv2 as a teacher model to train a 3D student backbone through pixel-to-point correspondence of features. While pre-training models on the target domain is a promising direction [45, 73], there is a lack of off-the-shelf models for autonomous driving, of a similar quality to DINOv2.
Pre-training 3D Backbones Self-supervised pre-training of 3D backbones started with single object scans (e.g. ShapeNet [13, 77]) similar to ImageNet in the case of images [17, 56, 95, 88, 60]. Several methods have been developed to pre-train on complex 3D indoor and outdoor scenes [81, 87, 74, 53]. Inspired by their vision counterpart, these methods also mainly use contrastive and reconstruction objectives. Some works leverage LiDAR specific augmentations [43] or use specialized tasks such as occupancy prediction consistent under sparse point sampling [8, 1] and different mixing strategies for pointclouds [52, 42, 79]. However, these methods are limited by the volume of diverse 3D pre-training data, owing to the difficulty in collecting such a large-scale dataset.
Cross-modal 2D-to-3D Knowledge Distillation: Recently, distilling knowledge from a pre-trained 2D image backbone (usually a VFM such as DINOv2 [54]) into a 3D LiDAR backbone has been shown to be an effective pre-training strategy. Initial works focused on single object pointclouds [22, 91] but recent works have demonstrated that this strategy is also suitable for complex scenes [50, 48]. SLidR [61] groups points and pixels into semantically coherent superpixels/superpoints based on the teacher vision model features. Other works use additional vision models such as SAM [37] or OpenSeeD [90] to extract pseudo-semantic regions for the same purpose [49, 83]. Recently, some works have also used spatial and temporal consistency to regularize learned representations [49, 83, 84, 94, 93]. Recently, ScaLR [57] demonstrated the potential of scaling up different distillation components such as the dataset, the 3D backbone and the 2D backbone while using a simple distillation loss that matches the corresponding point-pixel features directly. However, direct feature comparison impairs the 3D network’s ability to learn geometrically rich features. In this work, we propose an MLP projection head and the auxiliary spatial task of occupancy prediction, to guide the 3D network towards learning more spatially aware features.
Figure 2: Overview of the CleverDistiller framework. First, corresponding points and camera regions are identified based on sensor calibration data. Next, features from cameras and points are extracted using their respective backbones. Two tasks are employed for representation learning: First, the cross-modal KD loss is applied to distill features from the camera to the 3D backbone with our novel MLP projection head allowing the 3D network to learn complex semantics throughout the projection. Second, we employ an occupancy loss to learn spatially consistent feature representation. After pre-training, the 3D backbone can be utilized for various downstream tasks.
3 Method Description
In the following, we introduce our considered problem setting as well as our CleverDistiller framework shown in Fig. 2.
3.1 Problem Definition and Motivation
Given a point cloud $\mathcal{P}=\{\mathbf{p}_{i}\mid i=1,\ldots,N\}$ with $N$ points captured by a LiDAR sensor, where $\mathbf{p}_{i}\in\mathbb{R}^{4}$ denotes the coordinates of the point and the laser intensity. Let $\mathcal{I}_{t}=\{\mathbf{I}_{k}\mid k=1,\ldots,M\}$ denote $M$ surround camera images where $\mathbf{I}_{k}\in\mathbb{R}^{H\times W\times 3}$ denotes an image with height $H$ and width $W$. The objective of cross-modal distillation is to take adventage of a pre-trained vision foundation model $T_{\phi}(\cdot)$ to learn a LiDAR foundation model, $S_{\theta_{p}}(\cdot)$ by distilling the features from the image modality to the LiDAR modality.
Many recent works generate a set of class-agnostic or pseudo-semantic superpixels for each image using additional supervised or unsupervised segmentation and clustering models such as SAM [37] or OpenSeeD [90]. This introduces the implicit assumption of feature regularity within each mask or superpixel. Also, these masks and superpixels are not ground truths and require careful hyperparameter tuning to obtain suitable segments for a particular domain. This complicates the overall distillation process and we question the need for semantic priors in the presence of good features from a VFM.
Further, most of the previous works use a contrastive loss [49, 92, 83, 93, 84] which may create an unnecessary self-conflict issue. Specifically, in our approach we find corresponding point and image features, without relying on any additional processing steps or semantic priors. Assuming that the point cloud $\mathcal{P}_{t}$ and images $\mathcal{I}_{t}$ are calibrated, we find corresponding points and pixels by projecting the point cloud $\mathbf{p}_{i}=(x_{i},y_{i},z_{i})$ onto the image plane $(u_{i},v_{i})$ using the following sensor calibration parameters:
$\begin{bmatrix}u_{i}\\ v_{i}\\ 1\end{bmatrix}=\rho(i)=\frac{1}{z_{i}}\times\Gamma_{K}\times\Gamma_{c\leftarrow l}\times\begin{bmatrix}x_{i}\\ y_{i}\\ z_{i}\end{bmatrix},$
where $\Gamma_{K}$ denotes the camera-intrinsic matrix and $\Gamma_{c\leftarrow l}$ is the transformation matrix from LiDAR sensors to surround-view cameras. This enables a direct mapping between points and pixels in the image.
3.2 Cross-modal Learning Objectives
Let the student model, $S_{\theta_{p}}:\mathbb{R}^{N\times 4}\rightarrow\mathbb{R}^{N\times D_{p}}$ be a 3D backbone with trainable parameters $\theta_{p}$, which takes LiDAR points as input and outputs $D$-dimensional point features. Let $T_{\phi}:\mathbb{R}^{H\times W\times 3}\rightarrow\mathbb{R}^{\frac{H}{S}\times\frac{W}{S}\times D_{v}}$ be an image backbone with pre-trained parameters $\phi$ that takes images as input and outputs $D_{v}$-dimensional image features.
The objective function for such a self-supervised knowledge distillation is formulated as, $\mathcal{L}_{\text{distillation}}(\boldsymbol{q}_{p},\boldsymbol{q}_{v})$, where the primary objective is to transfer knowledge from the trained image backbone $T_{\phi}$ to the 3D backbone $S_{\theta_{p}}$. The point features are obtained as $\boldsymbol{q}_{p}=\mathcal{H}_{p}(S_{\theta_{p}}(\mathbf{p}))$ and image features $\boldsymbol{q}_{v}=T_{\phi}(\mathbf{I})$, where $\mathcal{H}_{p}:\mathbb{R}^{D_{p}}\rightarrow\mathbb{R}^{D_{v}}$ is a projection head that transforms the point features to the same dimensions as the VFM teacher. Inspired by ScaLR [57], we use a simple pointwise cosine similarity loss given by,
$\mathcal{L}_{\text{distillation}}(\boldsymbol{q}_{p},\boldsymbol{q}_{v})=\frac{1}{N}\sum_{i\in\{1,\ldots,N\}}\|\boldsymbol{q}_{p,i}-\boldsymbol{q}_{v,\rho(i)}\|_{2}.$ (1)
Such a loss is preferable to a contrastive loss as it avoids the complexities of careful negative sample selection, training sensitivity to batch composition and self-conflict issues while maintaining robust feature alignment.
3.3 Improving the Projection Head
Despite demonstrating that a simple approach can work for distillation, ScaLR yields subpar results compared to methods incorporating additional models, processing steps, or semantic priors [49, 83]. We found that while leveraging ScaLR approach can be beneficial, improving the projection head, a component overlooked in prior works, can significantly improve knowledge transfer.
We investigate the role of the projection head by evaluating the informativeness of the 3D backbone features using the RankMe metric [27]. The RankMe metric has been shown to be indicative of the informativeness of the learned representations and the downstream performance of self-supervised models. We find this to be valid for distillation features as well - features with a higher RankMe metric produce better downstream performance (see Fig. 3). ScaLR uses a linear 3D projection head and we discover that this results in less informative 3D backbone features (see Tab. 1). Using an MLP projection head improves the informativeness of the 3D backbone features as evidenced by the improved RankMe metric. Consequently, this results in significant downstream performance improvements (see Tab. 1). A deeper MLP with larger hidden dimensions results in the most informative backbone features and best downstream performance. A complex MLP projection head adds only a minimal increase in computational cost. Based on this analysis, we use a 3-layer MLP projection head with a similar formulation as the MLP head used while pre-training the DINOv2 model. Following these experiments, we set for using MLP projection head with 3 layers and 2048 hidden dim, that (as shown in Tabs. 1 and 3) yields the best results and retain the most informativeness without significant increase computational cost or memory.
3.4 Auxiliary Spatial Occupancy Task
Cross-modal distillation alone is insufficient for optimal performance as it ignores the rich spatial information unique to LiDAR that can be useful for downstream tasks. To enhance the LiDAR model’s spatial understanding and to encourage the learning of LiDAR specific features, we incorporate an auxiliary spatial task. While many works prioritize segmentation tasks where semantic knowledge is dominant, tasks such as 3D object detection (3DOD) require a deeper comprehension of spatial relationships. We experimented with auxiliary tasks like scene flow estimation and occupancy prediction, ultimately selecting occupancy as the most effective. Additionally, while previous works, such as SuperFlow [83], claim to enforce temporal consistency, our analysis revealed that their approach relies on semantic consistency (regions are matched based on the OpenSeeD class vocabulary rather than scene flow) across frames rather than on true temporal feature alignment.
Figure 3: Analysis of 1% finetuning MinkUnet [19] performance on nuScenes dataset (mIoU) vs RankMe metric of the 3D backbone features. We observe clear correlation between RankMe metric and 3D segmentation performance. The colors refers to the used distillation setup, $\bullet$ are using ViT-S/14 as a teacher and $\times$ ViT-B/14.
MLP Layers. Hidden dim. RankMe nuScenes KITTI Waymo
LP 1% 1% 1%
1 (Linear) - 1.89 40.76 53.01 44.42 47.98
2 2048 2.94 45.46 54.38 50.09 48.19
3 256 4.18 44.79 54.80 50.10 48.93
3 2048 4.24 46.36 55.01 50.15 49.08
Table 1: Comparison of different MLP architectures with the PT-v3 [75] 3D backbone model. The MLP projection improves performance as well as the RankMe metric. All scores in mIoU.
We incorporate an auxiliary spatial understanding task, specifically, occupancy prediction. This injects spatial and geometric information of the objects into the learned features, which we expect to be useful for tasks such as 3D object detection. The occupancy prediction loss is formulated based on ALSO [8]. We observe that incorporating the occupancy task improves the informativeness of the backbone features by enabling it to learn additional information complementary to those distilled from the VFM features (see improvement on RankMe metric when adding the occupancy task in Tab. 3). We find that this further boosts downstream performance in various tasks. The effective semantic information from the 2D backbone combined with the spatial knowledge encouraged by this task enable the model to understand and interpret complex 3D environments better. This also makes the model more robust to domain shifts and LiDAR corruption.
Following ALSO [8], for each 3D point $p$ sampled on the surface, we create two query points: $q_{\text{front}}=p-\delta p$ and $q_{\text{behind}}=p+\delta p$, where $\delta p$ is a small distance. We consider $q_{\text{front}}$ to be empty and $q_{\text{behind}}$ to be filled. Additionally, a third empty query point $q_{\text{sight}}$ is randomly picked in the segment between the sensor and $p$. An occupancy decoder head takes a concatenation of the 3D backbone features and the 3D coordinates of the associated point to predict the occupancy of that point and the intensity of that point. The occupancy loss consists of a binary cross-entropy loss $\mathcal{L}_{\mathrm{O}}$ based on the occupancy classification and a reconstruction MSE loss $\mathcal{L}_{\mathrm{Int}}$ based on the intensity prediction and use $\lambda=1$ similar to ALSO:
$\mathcal{L}_{\mathrm{occ}}=\mathcal{L}_{\mathrm{O}}+\lambda\mathcal{L}_{\mathrm{Int}}$ (2)
Overall, we combine two simple but significant design choices. First, we use an MLP projection head and cosine similarity loss. Second, we add a complementary auxiliary spatial task. Combined, these improve the informativeness and spatial consistency of the 3D backbone features as well as downstream performance. This combination of design choices leads to a more robust and efficient cross-modal distillation framework.
4 Results
4.1 Experiments setup
Implementation:
CleverDistiller is implemented using the Pointcept [20] codebase. Consistent with prior works we employ MinkUNet [19] as the 3D backbone and DINOv2 [54] (with ViT backbones [23]) as the 2D backbone, distilling from three variants: small (S), base (B) and large (L). Additionally, we experiment with a more flexible SOTA 3D backbone, PointTransformer-V3 (PT-v3) [75], to investigate if our method scales to larger models. For PT-v3, we use the default configuration and only increase the decoder dimensions to 256. The framework is pre-trained end-to-end on nuScenes using 4 GPUs for 50 and 100 epochs. We follow the same procedure to pre-train the baseline ScaLR model. We trained our model and ScaLR (official ScaLR results on MinkUnet are very limited [57]), while other results are based on what authors reported [51, 61, 49, 83, 84].
The distilled LiDAR backbone is then evaluated on a range of downstream tasks on various datasets. Firstly, we conduct segmentation experiments on nuScenes using linear probing and limited data finetuning scenarios using 4 GPUs for 100 epochs, utilizing the AdamW optimizer and OneCycle scheduler. The domain generalization study using Robo3D adheres to the same protocol as previous works. We also test our features for the 3D object detection (3DOD) task on the KITTI and Waymo datasets. For 3DOD, we use weights obtained during pre-training on nuScenes. We substitute PointRCNN [62] backbone (PointNet) with MinkUNet backbone (following SLiDR approach). Finally, we train for 80 epochs on partial data on Waymo and Kitti datasets. Our framework can also work with any other point-based 3D backbone.
Though most cross-modal image-to-LiDAR distillation works claim to pre-train for only 50 epochs, some recent works such as SuperFlow [83] utilize 3 LiDAR sweeps per sample, increasing the effective length and compute cost of the training by a factor of three (effectively 150 epochs). We characterize this using effective epochs in Tab. 4. For fairer comparison, we show results of our method for both 50 and 100 epochs of pre-training. CleverDistiller50 provides a direct comparison to previous methods that used only one LiDAR sweep per sample while CleverDistiller100 serves as a more efficient comparison to Superflow. After pre-training we fine-tune the models for 100 epochs.
We experiment on a total of nine datasets in our experiments, namely: nuScenes [9], SemanticKITTI [7] (SKITTI), KITTI [28], Waymo [64], ScribbleKITTI [69], RELLIS-3D [35], SemanticSTF [80], and DAPS-3D [39] and nuscenes-C from Robo3D [40].
Occupancy loss weight nuScenes SKITTI Waymo
LP 1% 1% 1%
0.0 55.72 61.79 56.30 60.78
0.01 56.49 61.40 57.01 62.14
0.05 58.49 62.54 62.19 61.95
0.2 57.39 61.65 58.96 62.39
1.0 53.89 58.99 56.78 62.30
Table 2: Comparison of different occupancy loss weights. All results in mIoU.
# MLP Occ Matrix Rank RankMe nuScenes SKITTI Waymo
LP 1% 1% 1%
(a) 22 1.89 40.76 53.01 44.42 47.98
(b) 25 2.33 41.67 52.16 48.19 48.11
(c) 33 4.24 46.36 55.01 50.15 49.08
(d) 41 5.27 49.81 56.90 50.59 50.99
Table 3: Ablation study of our components, showing the effectiveness of both the MLP projection head and the auxiliary occupancy task (Occ) indivisually. All results in mIoU.
Method Venue Distill Semantic Priors Effective Epochs nuScenes SKITTI Waymo
LP 1% 5% 10% 25% Full 1% 1%
Random - - - 8.10 30.30 47.84 56.15 65.48 74.66 39.50 39.41
PPKT [51] arXiv’21 ViT-S 50 38.60 40.60 52.06 59.99 65.76 73.97 43.25 47.44
SLidR [61] CVPR’22 ViT-S 50 44.70 41.16 53.65 61.47 66.71 74.20 44.67 47.57
Seal [49] NeurIPS’23 ViT-S 50 45.16 44.27 55.13 62.46 67.64 75.58 46.51 48.67
CleverDistiller50 - ViT-S 50 46.02 53.67 62.34 64.04 68.06 77.44 49.58 50.45
SuperFlow [83] ECCV’24 ViT-S 150 46.44 47.81 59.44 64.47 69.20 76.54 47.97 49.94
ScaLR [57] CVPR’24 ViT-S 100 40.76 53.01 61.58 63.47 67.56 74.51 44.42 47.98
LiMoE [84] CVPR’25 ViT-S 150 48.20 49.60 60.54 65.65 71.39 77.27 49.53 51.42
CleverDistiller100 - ViT-S 100 49.81 56.90 64.55 65.92 70.11 77.61 50.59 50.99
PPKT [51] arXiv’21 ViT-B 50 39.95 40.91 53.21 60.87 66.22 74.07 44.09 47.57
SLidR [61] CVPR’22 ViT-B 50 45.35 41.64 55.83 62.68 67.61 74.98 45.50 48.32
Seal [49] NeurIPS’23 ViT-B 50 46.59 45.98 57.15 62.79 68.18 75.41 47.24 48.91
CleverDistiller50 - ViT-B 50 48.40 55.80 63.49 65.56 69.06 78.17 49.83 52.80
SuperFlow [83] ECCV’24 ViT-B 150 47.66 48.09 59.66 64.52 69.79 76.57 48.40 50.20
ScaLR [57] CVPR’24 ViT-B 100 41.80 55.83 63.46 65.24 68.70 74.76 45.59 49.60
LiMoE [84] CVPR’25 ViT-B 150 49.07 50.23 61.51 66.17 71.56 77.81 50.30 51.77
CleverDistiller100 - ViT-B 100 51.89 59.80 66.44 67.65 69.53 78.49 51.48 53.56
PPKT [51] arXiv’21 ViT-L 50 41.57 42.05 55.75 61.26 66.88 74.33 45.87 47.82
SLidR [61] CVPR’22 ViT-L 50 45.70 42.77 57.45 63.20 68.13 75.51 47.01 48.60
Seal [49] NeurIPS’23 ViT-L 50 46.81 46.27 58.14 63.27 68.67 75.66 47.55 50.02
CleverDistiller50 - ViT-L 50 48.32 56.70 64.51 66.25 69.33 78.68 49.86 53.32
SuperFlow [83] ECCV’24 ViT-L 150 48.01 49.95 60.72 65.09 70.01 77.19 49.07 50.67
ScaLR [57] CVPR’24 ViT-L 100 40.12 55.78 63.28 64.76 68.19 75.09 44.85 50.34
LiMoE [84] CVPR’25 ViT-L 150 49.35 51.41 62.07 66.64 71.59 77.85 50.69 51.93
CleverDistiller100 - ViT-L 100 52.45 60.64 67.03 67.29 70.45 78.29 52.28 54.83
Table 4: Comparison of SOTA pretraining methods pretrained in self-supervised manner on nuScenes and fine-tuned on nuScenes, SemanticKITTI, and Waymo. All methods use MinkUNet-34 [19] as the backbone for the 3D semantic segmentation task. We also show which 2D VFM backbone was used as a teacher and if the approach uses pseudo semantic priors from off-the shelf segmentation models. LP denotes linear probing with a frozen backbone. All scores in mIoU. For LiMoE [84], we use LiMoE+Superflow.
Method Distill nuScenes SKITTI Waymo
LP 1% 5% 10% 25% 1% 1%
ScaLR [57] ViT-S 44.99 62.07 69.10 72.91 75.69 56.92 60.09
CleverDistiller ViT-S 58.49 62.54 71.09 73.89 76.64 61.95 62.19
ScaLR [57] ViT-B 41.62 63.54 71.74 73.84 76.56 60.74 57.17
CleverDistiller ViT-B 60.54 63.93 72.27 74.11 76.93 63.31 62.20
ScaLR [57] ViT-L 45.54 66.17 72.34 74.96 77.56 58.48 61.47
CleverDistiller ViT-L 61.92 66.18 73.13 75.18 77.57 62.23 65.94
Table 5: Comparisons of state-of-the-art pretraining methods pretrained on nuScenes and fine-tuned on SemanticKITTI and Waymo with specified data portions using PointTransformerV3 [76] as the 3D semantic segmentation backbone. LP denotes linear probing with a frozen backbone. All scores in mIoU.
Figure 4: Qualitative segmentation results comparing our CleverDistiller (right) to ScaLR (middle) and the ground truth (left). Legend corresponds to the ground truth. For ScaLR [57] and CleverDistiller we show prediction errors (in red), showcasing that our approach exhibits much less errors. Each row is a different sequence randomly sampled from nuScenes [9].
Evaluation Protocols:
Following standard conventions, we report the Intersection-over-Union (IoU) on each semantic class and mean IoU (mIoU) over all classes for downstream tasks. For 3D robustness evaluations, we follow Robo3D [40] and report the mean Corruption Error (mCE) and mean Resilience Rate (mRR). Evaluations are without test-time augmentations.
4.2 Ablation studies
We conduct ablation studies for the two components that we added: the MLP projection head and the auxiliary occupancy task. We pre-train a MinkUNet-34 model for 100 epochs using the ViT-S/14 variant of DINOv2. We use the 3-layer MLP head selected in Sec. 3.3 and experimented with different weights for the occupancy loss. From Tab. 2, we see that a weight of 0.05 is sufficient to incorporate additional spatial information. Then, we ablate the best versions of the two components Tab. 3. We observe that the addition of the MLP (c) improves the informativeness of the backbone features and provides a significant performance boost compared to the baseline (a) in Tab. 3. The occupancy task provides an orthogonal improvement that further improves the backbone features.
4.3 Performance Comparison
Leveraging data representation learning can significantly reduce the dependency on data annotations. We evaluate CleverDistiller against prior SOTA methods on three prominent datasets: nuScenes, SemanticKITTI, and Waymo, under scenarios with limited annotations for few-shot fine-tuning. As illustrated in Tab. 4 and Tab. 5, CleverDistiller demonstrates SOTA performance across all fine-tuning tasks as well as on linear probing on almost every scenario on both MinkUnet [19] and PTV3 [75]. Especially when fine-tuning on low data amounts on nuScenes we observe strong improvements of up to 10% mIoU, e.g. for 1% nuScenes data and ViT-L DINOv2 teacher. This performance indicates strong generalization of our learned feature representations, making our work highly relevant for use cases with low data availability. The tables also show that models pretrained through representation learning consistently surpass those initialized randomly, highlighting the benefits of data pretraining. Additionally, distillations from larger 2D networks yield consistent enhancements which is consistent with observations in previous works.
We can also observe in Fig. 4 that our predictions have fewer errors along the segmentation boundaries than those from ScaLR. Types of segmentation errors are not reflected by mIoU scores, but well-defined boundaries are crucial for the output quality. We provide additional qualitative of features and class-specific analysis in supplementary Sec. A2 and Sec. A3.
Method ScriKITTI Rellis-3D SemSTF DAPS-3D
1% 10% 1% 10% 50% 100% 50% 100%
Random 23.81 47.60 38.46 53.60 48.03 48.15 74.32 79.38
PPKT50 36.50 51.67 49.71 54.33 50.92 54.69 78.90 84.00
SLidR50 39.60 50.45 49.75 54.57 52.01 54.35 81.00 85.40
Seal50 40.64 52.77 51.09 55.03 53.46 55.36 81.88 85.90
Ours50 40.77 53.31 56.23 56.60 53.77 54.72 82.58 86.15
ScaLR100 36.45 49.16 47.91 48.86 52.10 54.40 - -
SF150 42.70 54.00 52.83 55.71 54.72 56.57 82.43 86.21
LiMOe150 43.95 55.96 53.74 56.67 55.60 57.31 83.24 86.68
Ours100 44.03 56.70 58.35 60.92 53.99 55.66 83.06 87.95
Table 6: Domain generalization study of various pretraining methods on the nuScenes, followed by fine-tuning on different 3D semantic segmentation datasets with specified data portions. All scores in mIoU.
Method Backbone mCE $\downarrow$ mRR $\uparrow$ mIoU $\uparrow$
Fog Rain Snow Blur Beam Cross Echo Sensor Avg Full
Full Random MinkU-34 112.20 72.57 62.96 70.65 55.48 51.71 62.01 31.56 59.64 39.41 54.18
PPKT50 [51] MinkU-34 105.64 75.87 64.01 72.18 59.08 57.17 63.88 36.34 60.59 39.57 56.60
SLidR50 [61] MinkU-34 106.08 75.99 65.41 72.31 56.01 56.07 62.87 41.94 61.16 38.90 56.83
Seal50 [49] MinkU-34 92.63 83.08 72.66 74.31 66.22 66.14 65.96 57.44 59.87 39.85 62.81
CleverDistiller50 MinkU-34 94.90 84.12 72.73 73.33 67.35 56.42 64.14 63.76 60.65 47.27 63.21
SuperFlow150 [83] MinkU-34 91.67 83.17 70.32 75.77 65.41 61.05 68.09 60.02 58.36 50.41 63.68
Scalr100 [57] MinkU-34 99.86 81.35 69.23 72.47 60.67 55.11 64.07 56.93 60.34 46.11 60.62
LiMoE150 MinkU-34 88.43 83.28 71.10 75.92 65.66 63.86 68.52 60.78 61.91 50.66 64.80
CleverDistiller100 MinkU-34 91.28 87.42 72.83 70.91 66.03 60.18 66.85 68.04 60.17 53.63 64.83
LP PPKT50 [51] MinkU-34 183.44 78.15 30.65 35.42 28.12 29.21 32.82 19.52 28.01 20.71 28.06
SLidR50 [61] MinkU-34 179.38 77.18 34.88 38.09 32.64 26.44 33.73 20.81 31.54 21.44 29.95
Seal50 [49] MinkU-34 166.18 75.38 37.33 42.77 29.93 37.73 40.32 20.31 37.73 24.94 33.88
CleverDistiller50 MinkU-34 157.50 84.84 42.45 43.45 38.77 37.85 41.01 43.11 39.17 26.55 39.05
SuperFlow150 [83] MinkU-34 161.78 75.52 37.59 43.42 37.60 39.57 41.40 23.64 38.03 26.69 35.99
Scalr100 [57] MinkU-34 173.18 78.91 37.55 37.96 36.29 33.64 33.06 23.01 33.62 23.70 33.59
LiMoE150 MinkU-34 155.77 78.23 40.35 45.28 39.14 42.10 44.21 27.33 39.20 29.49 38.39
CleverDistiller100 MinkU-34 151.21 89.99 43.96 46.91 41.20 41.05 42.15 45.67 41.30 28.85 41.39
Table 7: 3D robustness study of state-of-the-art pretraining methods under corruption and sensor failure scenarios in the nuScenes-C dataset from the Robo3D benchmark [40]. Full denotes fine-tuning with 100% labeled data. LP denotes linear probing with a frozen backbone. All mCE (↓), mRR (↑), and mIoU (↑) scores are given in percentage (%). Best scores in each configuration are shaded with colors.
4.4 Domain generalization
We test our method, pre-trained on nuScenes, on the ScriKITTI, Rellis-3D, SemSTF, and DAPS-3D datasets to show its domain generalization capabilities, as shown in Tab. 6. Each dataset presents unique challenges, and our method consistently performs well. Our approach consistently achieved the highest scores in most cases compared to other SOTA pretraining methods. For instance, on the ScriKITTI dataset, our method reached 40.77 mIoU with 1% data and 53.31 with 10% data, outperforming all other methods. Similarly, on the Rellis-3D dataset, our method achieved 56.23 mIoU with 1% data and 56.60 with 10% data, again leading the performance charts. Even on the SemSTF and DAPS-3D datasets, our method demonstrated strong results, particularly excelling with 82.58 mIoU at 50% data and 86.15 at 100% data on DAPS-3D. These results underscore the robustness and generalization of our approach in 3D semantic segmentation tasks.
4.5 Robustness
The robustness of 3D perception models against unprecedented conditions directly correlates with the model’s applicability to real-world applications. We compare our method with prior models in the nuScenes-C dataset from the Robo3D [40] benchmark and show results in Tab. 7. Our method demonstrates superior performance, particularly in corrupted scenarios where other methods struggle, such as when the LiDAR data is much sparser.
CleverDistiller shows improved robustness over random initialization and other methods, achieving the lowest mean Corruption Error (mCE) for linear probing and nearly the lowest for full pre-training. Additionally, our method achieves the highest mean relative robustness (mRR) in all trials, indicating strong resistance to corruptions. For example, in the Fog scenario, our method achieves an mIoU of 72.83, compared to 62.96 for random initialization. The most notable result is in the Cross corruption scenario, which involves sparse LiDAR data. While other methods experience a performance drop of approximately 50% on LP and 10-20% on full, CleverDistiller maintains high performance, showcasing its adaptability to sparse data.
4.6 3D Object Detection
Method KITTI Waymo
5% 10% 20% 5% 10% 20%
Random init. 56.1 59.1 61.6 18.0 21.9 28.5
PPKT50 57.8 60.1 61.2 - - -
SLidR50 57.8 61.4 62.4 - - -
CleverDistiller50 58.3 63.7 66.8 22.7 24.6 26.4
ScaLR100 56.2 62.3 66.5 20.1 28.1 28.2
CleverDistiller100 59.8 66.6 67.1 22.2 28.7 29.6
Table 8: Evaluation of our pre-training approach on 3D object detection benchmarks. All results in mAP.
For 3DOD, we use the same features distilled in SSL as we do for segmentation. This means we have one model, one distillation run, and the same features for multiple tasks, ensuring efficiency and consistency. 3D backbone is pretrained in an unsupervised manner and then used as a backbone in the 3DOD task. Unlike other methods such as Olivine [92] and CMCR [93], which need to distill features to a different network architecture like VoxelNet, our approach uses MinkUNet, making the features useful for both segmentation and 3DOD tasks. In our experiments shown in Tab. 8, we benchmarked our method on the KITTI and Waymo datasets. Our method consistently outperformed the previous best method by approximately 5-10%, showing that the features are versatile and effective across multiple tasks.
5 Conclusions
We introduce CleverDistiller, a framework for cross-modal distillation that achieves SOTA performance using simple and spatially consistent cross-modal distillation. We show that the projection head design plays a crucial role in improving informativeness of the learned representations. Crucially, we observe that our MLP projection head enables better semantic information flow from the 2D VFM model to the 3D backbone. Incorporating an auxiliary spatial tasks, such as occupancy prediction, provides additional supervision signals that enhance the spatial reasoning of the model. The results, both qualitative and quantitative, demonstrate that the learned features are an excellent basis for various downstream tasks, especially in sparse data scenarios where we see improvements by up to 10% mIoU. Since our proposed improvements are widely applicable to other architectures we believe that our contributions will have a high adoption rate by future work on cross-modal knowledge distillation.
Supplementary Material
A1 Additional Experiments
We use the Mix3D [52] augmentation during pre-training and show an ablation study for this choice in Tab. 9. We find that Mix3D provides a modest performance boost. For pre-training both ScaLR and our CleverDistiller we use the Mix3D augmentation.
Mix3D nuScenes KITTI Waymo
LP 1% 1% 1%
48.95 55.36 49.03 50.09
49.81 56.90 50.59 50.99
Table 9: Ablation study of Mix3D during distillation
A2 Class-wise results
In Tab. 10 we compare the class-wise performance of CleverDistiller with other methods after finetuning on 1% of the training data. We generally outperform other methods, including the most recent LiMoE on most classes Tab. 10.
Method mIoU
barrier
bicycle
bus
car
construction vehicle
motorcycle
pedestrian
traffic cone
trailer
truck
driveable surface
other flat
sidewalk
terrain
manmade
vegetation
ViT-S
SLidR 41.2 0.0 0.0 26.6 72.0 12.4 15.8 51.4 22.9 11.7 35.3 92.9 36.3 58.7 63.6 81.2 82.3
Seal 44.3 20.0 0.0 19.4 74.7 10.6 45.7 60.3 29.2 17.4 38.1 93.2 26.0 58.8 64.5 81.9 81.9
CleverDistiller50505050 53.7 59.1 0.0 61.4 83.4 1.2 40.1 62.1 23.1 28.0 56.3 93.7 53.3 62.2 68.8 83.5 82.6
ScaLR 53.0 58.4 0.0 64.9 83.3 10.5 20.7 59.6 24.6 26.4 55.8 93.9 51.2 63.4 68.5 83.7 83.1
SuperFlow 47.8 38.2 1.8 25.8 79.0 15.3 43.6 60.3 0.0 28.4 55.4 93.7 28.8 59.1 59.9 83.5 83.1
LiMoE 49.6 39.9 4.6 27.3 80.2 17.1 45.4 61.2 6.2 29.5 58.4 94.0 34.2 62.3 64.6 84.1 84.5
CleverDistiller100100100100 56.9 60.2 0.0 71.2 84.9 11.8 47.1 65.1 26.6 31.4 61.7 94.2 54.1 64.9 69.5 84.2 83.5
ViT-B
PPKT 40.9 0.0 0.0 24.5 73.5 12.2 7.0 51.0 13.5 15.4 36.3 93.1 40.4 59.2 63.5 81.7 82.2
SLidR 41.6 0.0 0.0 26.7 73.4 10.3 16.9 51.3 23.3 12.7 38.1 93.0 37.7 58.8 63.4 81.6 82.7
Seal 46.0 43.0 0.0 26.7 81.3 9.9 41.3 56.2 0.0 21.7 51.6 93.6 42.3 62.8 64.7 82.6 82.7
CleverDistiler50 55.8 60.7 0.0 68.8 85.1 22.0 28.9 61.8 27.6 30.1 61.0 94.0 52.9 63.7 69.1 84.0 82.9
SuperFlow 48.1 39.1 0.9 30.0 80.7 10.3 47.1 59.5 5.1 27.6 55.4 93.7 29.1 61.1 63.5 82.7 83.6
ScaLR 55.8 60.8 0.0 69.7 85.1 20.4 30.6 60.8 27.8 29.5 58.1 94.2 55.6 64.3 69.3 84.1 83.1
LiMoE 50.2 41.5 3.8 32.2 81.7 12.9 49.3 61.1 7.3 29.3 57.8 94.2 35.1 62.9 65.4 84.0 84.8
CleverDistiller 100 59.8 61.8 0.0 72.8 85.7 35.2 49.6 65.4 33.6 32.9 62.6 94.5 59.5 65.6 69.5 84.7 83.5
ViT-L
PPKT 42.1 0.0 0.0 24.4 78.8 15.1 9.2 54.2 14.3 12.9 39.1 92.9 37.8 59.8 64.9 82.3 83.6
SLidR 42.8 0.0 0.0 23.9 78.8 15.2 20.9 55.0 28.0 17.4 41.4 92.2 41.2 58.0 64.0 81.8 82.7
Seal 46.3 41.8 0.0 23.8 81.4 17.7 46.3 58.6 0.0 23.4 54.7 93.8 41.4 62.5 65.0 83.8 83.8
CleverDistiller 50 56.7 63.4 0.0 69.7 85.4 19.3 29.4 62.1 30.4 31.9 62.4 94.3 56.2 65.3 70.0 84.1 83.0
SuperFlow 50.0 44.5 0.9 22.4 80.8 17.1 50.2 60.9 21.0 25.1 55.1 93.9 35.8 61.5 62.6 83.7 83.7
LiMoE 51.4 45.3 4.1 25.3 82.2 18.4 52.5 61.8 22.3 26.4 56.2 94.3 37.6 63.3 63.9 84.4 85.0
ScaLR 55.8 65.0 0.0 68.7 85.3 15.9 27.7 61.4 28.3 31.0 61.8 94.2 51.6 64.6 70.0 84.1 83.0
CleverDistiller 100 60.6 65.5 0.0 72.3 88.3 37.2 37.4 65.1 35.3 40.9 71.9 94.5 55.8 66.5 70.6 85.0 83.8
Table 10: Per-class IoU scores of state-of-the-art pretraining methods pretrained and fine-tuned on nuScenes with 1% annotations. All IoU scores are given in percentage (%).
A3 Qualitative analysis of features
Similarity maps presented in Fig. 5 illustrate the segmentation ability of our pretrained model, CleverDistiller. The query points include “truck”, “car”, “driveable surface”, and “vegetation”. CleverDistiller shows strong semantic discriminative ability without fine-tuning, similar to features distilled from camera foundational models. This effectiveness can be attributed to three key aspects: 1) Cosine similarity loss, which ensures that features are aligned; 2) An MLP projection head that retains information during projections, unlike previous methods using linear layers, helping the features to become more aligned; 3) An additional 3D auxiliary spatial task that improves the 3D features and understanding of the 3D backbone. The visualization of feature similarities in Fig. 5 highlights how CleverDistiller’s features align with those of the teacher model. Rows 1 and 3 show similarities of query pixels to other pixels based on the teacher model’s image features, while rows 2 and 4 show similarities of query points to other points based on CleverDistiller’s distilled backbone features. Yellow denotes high similarity and blue denotes low similarity, with query points/pixels marked by red crosses. Best viewed in color.
Figure 5: Visualization of feature similarities of CleverDistiller. Rows 1 and 3 show similarities of query pixel to other pixels based on teacher model’s image features. Rows 2 and 4 show similarities of query point to other points based on the distilled backbone’s features. Best viewed in color. The query points / pixels are denoted using red crosses. Yellow denotes high similarity and blue denotes low similarity.